### v1

# SpecServe: Efficient and SLO-Aware Large Language Model Serving with Adaptive Speculative Decoding

Abstract
Large Language Model (LLM) services often face challenges in achieving low inference latency and meeting Service Level Objectives (SLOs) under dynamic request patterns. Speculative decoding, which exploits lightweight models for drafting and LLMs for verification, has emerged as a compelling technique to accelerate LLM inference. However, existing speculative decoding solutions often fail to adapt to varying workloads and system environments, resulting in performance variability and SLO violations. In this paper, we introduce SpecServe, an efficient LLM inference system that dynamically adjusts speculative strategies according to real-time request loads and system configurations. SpecServe proposes a theoretical model to understand and predict the efficiency of speculative decoding across diverse scenarios. Additionally, it implements intelligent drafting and verification algorithms to guarantee optimal performance while achieving high SLO attainment. Experimental results on real-world LLM traces demonstrate that SpecServe consistently meets SLOs and achieves substantial performance improvements, yielding 1.14$\times$-14.3$\times$ speedups over state-of-the-art speculative inference systems.
1 Introduction
Large Language Models (LLMs) have demonstrated their versatility across a broad spectrum of tasks [3], such as translation [24], question answering [16], and mathematical reasoning [7]. This success has led to the rapid adoption of LLM-based services, which accept users’ real-time queries, run LLM inference, and return generated responses. During LLM inference, the decoding phase generates output tokens sequentially (i.e., autoregressive decoding), which typically dominates end-to-end request latency. Hence LLM inference systems need to meet stringent performance requirements for token generation—often specified as Service-Level Objectives (SLOs) [40].
Speculative decoding has recently emerged as an appealing solution for accelerating LLM inference [18, 5]. It leverages a lightweight draft model to generate multiple draft tokens and then runs the target model (i.e., LLM) to verify them all at once. Compared with autoregressive decoding, this approach guarantees the same output results without needing to load LLMs for every token generation, effectively improving GPU utilization and reducing inference latency.
However, existing speculative decoding solutions often suffer from suboptimal performance when dealing with dynamic request patterns and varying system configurations, which are typical scenarios in real-world online LLM services. These solutions generally employ a pre-defined speculative length, i.e., the number of draft tokens generated in each round. While a larger speculative length increases the chance to accept more valid tokens in verification, it incurs higher latency and risks resource wastage due to potential token rejections [22]. Additionally, we observe that the optimal configuration of speculative length depends on the interplay between request loads, workload categories, and hardware resources (see §2.3). Consequently, using a pre-defined speculative length fails to consistently deliver satisfactory performance under changing workloads and environments.
Moreover, current speculative decoding solutions often do not take into account token-level SLOs of LLM inference, failing to ensure high SLO attainment. Although speculative decoding can achieve significantly lower inference latencies compared to autoregressive decoding, the uncertain token acceptance during verifications can lead to considerable fluctuations in overall token generation latency, such as Time Per Output Token (TPOT). This variability makes it challenging for speculative decoding to consistently meet SLO requirements, thereby affecting the ability to provide users with consistent and efficient services.
Therefore, we advocate that an efficient speculative inference system should enable fine-grained control over speculative lengths and beware of dynamic workloads and SLO requirements. Following this idea, we present SpecServe, a system that employs adaptive speculative decoding to reduce the end-to-end latency of online LLM inference. It dynamically determines the optimal speculative length for varying request patterns, system configurations, and SLO constraints. By accommodating these changing conditions, SpecServe strives to achieve lower inference latency and ensure SLO attainment. However, achieving adaptive speculative decoding in SpecServe presents three key challenges.
The first challenge is to build an efficiency model for speculative decoding that reflects dynamic system loads and diversified request categories. In online LLM inference, system load can be simply represented by metrics such as batch size and context length, but it is challenging to measure the impact of request complexity on the efficiency of speculative decoding. This is because it is not possible to know during the drafting phase whether the draft tokens will be accepted in the verification stage. Based on our observations, the acceptance rate of draft tokens is directly proportional to the confidence score at which they are generated by the draft model [8, 20]. Therefore, we propose the adaptive drafter, which integrates the efficiency model of speculative decoding into the execution of the draft model. It directly uses the confidence score of draft tokens to update their acceptance rate in real-time. Then, it employs a simple yet effective linear model to depict the impact of system load and system configuration on the execution time of speculative decoding. Ultimately, the optimal speculative length is determined collaboratively by the acceptance rate of draft tokens and the execution time of speculative decoding.
Secondly, SpecServe requires fine-grained control over the speculative length to deliver optimal inference performance. Due to the diversity of request categories, applying the same speculative length to different requests within the same batch still fails to achieve optimal latency. Therefore, we propose a succinct and efficient solution, named the confidence prior verifier. The essence of this solution lies in decoupling the drafting and verification phases, and leveraging the acceptance rate of draft tokens recorded by the adaptive drafter, as well as a speculative decoding execution time model. Through this approach, we can significantly reduce resource wastage caused by verifying useless draft tokens, by only allowing the target model to verify those tokens that are more likely to be accepted, without differentiating the speculative length of each request.
The third challenge is to improve SLO attainment for online LLM inference services. The SLO violations in speculative decoding originate from the discrepancy between the expected and actual number of accepted tokens, indicating that the additional decoding cost does not correspond to a proportional increase in generated tokens. To mitigate the issue of SLO violations, we leverage SLO constraints to control the speculative length. We integrate these SLO constraints directly into the efficiency model of speculative decoding, resulting in the development of a SLO-aware efficiency estimator for speculative decoding. This SLO-aware efficiency estimator considers not only the speed of token generation but also the quality of these tokens in meeting predefined SLO constraints.
We have implemented SpecServe atop vLLM [17] and evaluated it using real-world LLM inference traces. Evaluation results show that SpecServe can effectively adapt to varying workloads and system configurations, consistently achieving low inference latency and high SLO attainments. Compared with state-of-the-art speculative inference systems, SpecServe enables 1.14$\times$ to 14.3$\times$ speedups while meeting SLO requirements.
To summarize, we make the following contributions:
  • We proposed SpecServe, an efficient LLM inference system with adaptive speculative decoding.
  • We designed a fine-grained control mechanism for speculative length (i.e., adaptive drafter and confidence prior verifier), which can adapt to varying request patterns and system configurations and achieve optimal performance with theoretical guarantee.
  • We designed an efficient algorithm (i.e., SLO-aware efficiency estimator) to achieve high SLO attainment under varying performance requirements.
  • We evaluated SpecServe using real-world production traces, which demonstrates that SpecServe can significantly improve inference performance by up to 14.3 $\times$ compared with baselines.
2 Background and Motivation
2.1 LLM Inference Service
Figure 1: Different request patterns in real production traces.
In LLM inference [1, 2], the model accepts a list of tokens $\mathbf{x}=(x_{1},\ldots,x_{n})$, and outputs the conditional probability distribution of the next token $(x_{n+1}|x_{1},\ldots,x_{n})$. The trained LLMs are typically deployed onto inference systems to provide online services, where they receive requests consisting of token lists and generate a complete output sequence $(x_{n+1},\ldots,x_{n+T})$ by predicting the next token $T$ times.
In real-world LLM inference systems, they usually need to handle dynamic request arrivals from a variety of workloads. Fig. 1 illustrates two representative LLM inference traces from Azure [29]. The first pattern (trace 1) demonstrates sporadic bursts of requests, characterized by a generally low request rate but with extremely high peaks; the second pattern (trace 2) displays lower request volatility, yet still reflects the dynamic nature of system load. A key requirement of LLM inference is to meet stringent latency requirements (i.e., SLOs), which necessitates inference systems to efficiently handle dynamic requests and achieve low-latency token generation [10, 40, 41, 39].
Request batching is a common approach to improve inference performance and resource efficiency for LLM inference [38, 12]. It fuses multiple requests associated with the same model as a batch, and executes them together to fully exploit parallel computing resources of hardware accelerators such as GPUs. Mainstream LLM inference systems such as TensorRT-LLM [28] and vLLM [17] offer the ability of batching for enhanced efficiency. While request batching enhances GPU utilization during decoding, this method fails to address inefficiencies in scenarios with persistently low per-model request rates. According to real-world inference traces (Fig. 2), LLM services frequently experience prolonged periods of low request rates—a pattern consistent with prior observations of inference workloads[39]. These findings underscore a critical requirement for LLM inference systems: maintaining high computational efficiency even under sustained low request arrival rates.
2.2 Speculative Decoding
Figure 2: A single-step speculative decoding process with a speculative length of 1. We primarily discuss the greedy sampling, where the token with the highest confidence score is selected as the output instead of the whole distribution.
Speculative decoding addresses the inefficiency of traditional autoregressive decoding in LLM inference systems. This approach utilizes a small draft model to predict the following tokens and verify them at once, improving resource efficiency and end-to-end performance. As shown in Fig. 2, speculative decoding can be divided into two parts: drafting and verification. During the drafting phase, given a list of tokens $\mathbf{x}=(x_{1},\ldots,x_{n})$, speculative decoding uses an efficient draft model to generate a draft sequence $\mathbf{y}=(y_{1},\ldots,y_{SL})$, where $SL$ is the speculative length. In the verification process, the draft sequence $\mathbf{y}=(y_{1},\ldots,y_{SL})$ is input into the target model in parallel for a single forward pass, generating multiple output distributions $P(y_{1}|x_{1:n}),\ldots,P(y_{SL}|x_{1:n},y_{1:SL-1})$. Based on these output distributions and specific acceptance criteria, all tokens $(y_{1},\ldots,y_{m-1})$ before the rejected token $y_{m}$ that first fails the acceptance criteria will be selected as accepted tokens, plus a bonus token $z_{m}$ resampled from distribution $P(y_{m}|x_{1:n},y_{1:m-1})$, to form the output sequence $(y_{1},\ldots,y_{m-1},z_{m})$ generated by speculative decoding. Since the acceptance of tokens is still determined by the output distribution of the target model, speculative decoding is lossless compared to auto-regressive decoding.
The speculative length, by controlling the number of candidate tokens generated by the draft model, directly impacts both the latency benefits and resource consumption of speculative decoding. Current mainstream methods [18, 33, 26, 4, 19] rely on offline analysis to determine a static speculative length, with its optimal value typically identified through heuristic search on specific datasets and under fixed hardware configurations. Furthermore, while a few methods, such as SmartSpec [22], are capable of setting the speculative length for each decoding step, their coarse-grained control still fails to adapt to the fine-grained variations present in real-world request scenarios.
2.3 Limitations of Existing Solutions
Current speculative decoding approaches face three fundamental limitations in the practical deployment of LLM inference systems. These limitations stem from their lack of adaptability to dynamic request patterns, system configurations, and SLO constraints.
Figure 3: Speedup across various speculative lengths under diverse request quantities and types. The red dashed line represents the speedup value of 1.
Limitation #1: Inefficiency under dynamic workloads. As shown in Fig. 3, conventional speculative decoding methods, which rely on fixed speculative lengths, fail to adapt to dynamic changes in request patterns. Under high-load scenarios (batch size 128), excessively long speculative lengths lead to: (1) increased computational overhead during the verification phase due to redundant tokens generated by the draft model; (2) resource contention caused by batch sizes exceeding device memory. Conversely, under low-load scenarios (batch size 1), excessively short speculative lengths fail to fully utilize idle computational resources, leaving the acceleration potential of speculative decoding untapped. Workload characteristics represent another critical variable. We demonstrate this by analyzing workloads extracted from RAG [14], mathematical reasoning [7], and question-answering [16] datasets, each exhibiting different optimal speculative lengths. Evidently, applying the same speculative length across different requests results in both redundant token generation and wasted acceleration potential, making it difficult to achieve optimal latency optimization.
Figure 4: Speedup across different speculative lengths for various draft-target model pairs on diverse computing platforms.
Limitation #2: Agnostic to system configurations. As illustrated in Fig. 4, when hardware configurations or model pairings change, the fixed-length strategy of traditional speculative decoding fails to maintain optimality. This phenomenon arises from changes in the hardware roofline model: when the arithmetic intensity of the target model is below the platform’s balance point, parallel token processing during the verification phase can effectively improve computational efficiency [25]. Conversely, in compute-bound scenarios, the probabilistic nature of speculative decoding reduces its token generation efficiency compared to autoregressive decoding, resulting in latency degradation.
Limitation #3: SLO violations. As revealed in Fig. 5, while speculative decoding can achieve higher acceleration ratios compared to autoregressive decoding, its SLO attainment may be significantly lower. This issue arises from the probabilistic variability of speculative decoding’s acceleration effects: longer speculative lengths increase the expected acceleration ratio but also raise the risk of exceeding the TPOT limit (e.g., when the number of tokens actually generated is lower than expected, the TPOT for that speculative decoding attempt will significantly exceed its expected value). This uncertainty prevents speculative decoding from simultaneously meeting the high-performance and high-reliability requirements of inference systems.
2.4 Key Insight and Challenges
Figure 5: Relationship between SLO attainment and end-to-end latency speedup under different speculative lengths. The red dashed line represents the speedup value of 1 and the SLO attainment value of 0.9.
Key Insight: Fine-grained dynamic control. These limitations discussed in §2.3 reveal a core contradiction: the potential benefits of speculative decoding are constrained by its rigid execution framework. We propose a fundamental solution which transforms the speculative length from a preset parameter into a dynamically controlled variable. Specifically, the system needs to dynamically adjust the speculative length for each decoding step in real-time, based on workload characteristics, system configurations, and SLO constraints. This mechanism enables speculative decoding to both mitigate the risk of resource overload (e.g., by shortening speculative lengths under high-load conditions) and maximize computational resource utilization (e.g., by extending speculative lengths under low-load conditions), thereby overcoming the performance boundaries of static strategies.
However, achieving fine-grained dynamic control of speculative length faces three major challenges.
Challenge #1: How to accurately characterize the dynamic relationship between efficiency and speculative length? The relationship between the efficiency of speculative decoding and speculative length is influenced by dynamic request patterns and diverse system configurations in real-world scenarios (§2.3). To accurately characterize this relationship, we need to construct an effective and general-purpose efficiency model for speculative decoding. This model can guide the design of speculative strategies under varying scenarios for improved overall efficiency.
Challenge #2: How to timely and flexibly control the speculative length? The root of this challenge lies in two main factors: (1) Lag in efficiency estimation: The estimation of decoding efficiency depends on the actual execution results of the draft model. This dependency prevents the system from preemptively evaluating the validity of a given speculative length before generating the draft, leading to delay in controlling. (2) Rigid system architecture constraints: All requests within the same batch have the same number of draft model executions, which results in the default speculative length being the same for all requests, thereby neglecting fine-grained control.
Challenge #3: How to balance the efficiency of speculative decoding and SLO attainment? Due to the probabilistic fluctuations in the acceleration effects of speculative decoding, there is a conflict between maximizing the efficiency of speculative decoding and meeting SLO requirements (§2.3). Therefore, we need to incorporate SLO constraints and the risk of TPOT exceeding expectations into the efficiency model, allowing the system to balance acceleration gains and SLO violation risks in real-time when controlling speculative length.
Figure 6: An overview of SpecServe’s modules: adaptive drafter, confidence prior verifier, and SLO-aware efficiency estimator.
3 SpecServe’s Overview
We proposed SpecServe, an efficient LLM inference system with adaptive speculative decoding. To accurately characterize the dynamic relationship between efficiency and speculative length (Challenge #1), we establish an effective and general-purpose efficiency model for speculative decoding (§4). Based on this model, we implement three components to achieve fine-grained speculative decoding, as shown in Fig. 6.
Adaptive Drafter (§5.1). To address the lag in efficiency estimation (Challenge #2-1), SpecServe introduces an adaptive drafter, integrating efficiency estimation into the drafting phase to achieve proactive, step-level speculative length control. As illustrated in Fig. 6, before each execution of the draft model, the speculative controller uses historical confidence scores to estimate the potential efficiency gain of the current speculative length in real-time. After executing the draft model, the generated confidence scores are recorded and used to adjust the efficiency of the current length to ensure accurate speculative length control.
Confidence Prior Verifier (§5.2). To address the rigid system architecture constraints (Challenge #2-2), SpecServe employs a confidence prior verifier to prioritize the verification of tokens with high acceptance rates, thereby decoupling the drafting phase from the verification phase and enabling fine-grained request-level speculative length control. As shown in Fig. 6, before being input into the target model, the draft token set undergoes draft token elimination, where tokens are progressively removed in ascending order of acceptance rate until the draft token set with the highest efficiency is selected.
SLO-Aware Efficiency Estimator (§5.3). The key operations of SpecServe are based on the evaluation of the efficiency of speculative decoding. Similar to vLLM [17] and Orca [38], we use throughput to measure decoding efficiency and propose an efficiency estimator. This estimator employs the efficiency model of speculative decoding to adapt to dynamic request patterns and diverse system configurations. To achieve SLO awareness (Challenge #3), the efficiency estimator incorporates SLO constraints to support rapid trade-offs between the efficiency of speculative decoding and SLO compliance.
4 Modeling the Efficiency of Speculative Decoding
To evaluate the efficiency of speculative decoding, we adopted the same throughput metric as used in existing LLM inference systems [17, 38], which is defined as $Throughput=NAT/T_{sd}$. We developed two models: one for speculative decoding time ($T_{sd}$) and another for the number of accepted tokens ($NAT$). Based on these models, we investigated the relationship between the efficiency of speculative decoding and the speculative length, finding that the efficiency of speculative decoding initially increases and then decreases, or it monotonically decreases, with an increase in speculative length. This discovery provides a theoretical foundation for implementing efficient speculative length control.
4.1 The Time for Speculative Decoding
The time for speculative decoding $T_{sd}$ includes the execution time of the draft model ($T_{d}$) and the execution time of the target model ($T_{v}$), which can be expressed as:
$T_{sd}=T_{d}+T_{v}$ (1)
Modeling the forward propagation time. The execution of both the draft and the target models fundamentally involves multiple instances of forward propagation. Similar to previous work [22], we simplify the time for forward propagation as:
$T_{fwd}(N_{c},N_{b})=\alpha*N_{c}+\gamma*N_{b}+\delta$ (2)
, where $N_{c}$ and $N_{b}$ respectively represent the number of context tokens and batch tokens in the batch, and the performance coefficients $(\alpha,\gamma,\delta)$ depend on the computing platform, parallelism, and the model.
Modeling the execution time of draft and target models. The execution time of the draft model is the sum of the times for multiple forward passes, which can be expressed as:
$T_{d}$ $=\sum_{i=1}^{SL}\big{(}\alpha_{d}*N_{dc}(i)+\gamma_{d}*N_{db}(i)+\delta_{d}\big{)}$ (3)
$=\alpha_{d}*\sum_{i=1}^{SL}N_{dc}(i)+\gamma_{d}*\sum_{i=1}^{SL}N_{db}(i)+\delta_{d}*SL$ (4)
, where $(\alpha_{d},\gamma_{d},\delta_{d})$ is the performance coefficients of the draft model on the given computing platform, and $N_{dc}(i),N_{db}(i)$ are the number of context tokens and batch tokens in the batch during the $i^{th}$ forward propagation of the draft model.
Similarly, the execution time of the target model can be represented as the time for one forward propagation to process all speculative inputs, as shown in Eq. 5. We can relate the number of context tokens $N_{vc}$ and batch tokens $N_{vb}$ in each batch of the target model’s forward propagation to $N_{dc}$ and $N_{db}$, as shown in Eq. 6.
$T_{v}$ $=\alpha_{v}*N_{vc}+\gamma_{v}*N_{vb}+\delta_{v}$ (5)
$=\alpha_{v}*\sum_{i=1}^{SL+1}N_{dc}(i)+\gamma_{v}*\sum_{i=1}^{SL+1}N_{db}(i)+\delta_{v}$ (6)
, where $(\alpha_{v},\gamma_{v},\delta_{v})$ is the performance coefficients of the target model on the given computing platform.
Figure 7: Average acceptance rates for different confidence score intervals of the draft model. The red dashed line connects (0,0) and (1,1) to aid in visual assessment.
4.2 The Number of Accepted Tokens
Given the dynamic system loads and diverse request categories, it is impractical to estimate the number of accepted tokens for any given request using overly complex conditions. Therefore, we propose an acceptance token number model that does not rely on any conditions:
$NAT=\sum_{i=1}^{BS}(\sum_{j=1}^{SL}AR_{i,j}+1)=\sum_{i=1}^{BS}(\sum_{j=0}^{SL}AR_{i,j})$ (7)
, where $AR_{i,j}$ denotes the probability that the $j^{th}$ draft token of the $i^{th}$ request is accepted in the current step. Since every request has a bonus token that is inevitably accepted, the sum of the acceptance probabilities must also include this bonus token. In fact, autoregressive decoding can be considered as speculative decoding with a speculative length $SL=0$, accepting only one bonus token. To unify autoregressive decoding with speculative decoding, we define $AR_{i,0}=1$.
Figure 8: The relationship between estimated and actual number of accepted token with speculative length.
Modeling acceptance rate. The experimental results shown in Fig. 7 demonstrate that the acceptance rate of draft tokens under different combinations of draft and target models is positively correlated with their confidence scores in the draft model. Similar to previous works [8, 20], this finding suggests that we can model acceptance rates using confidence levels.
Considering the sequential dependency among tokens, we represent the probability that the $j^{th}$ draft token of the $i^{th}$ request is accepted in the current speculative decoding step as the probability that all previous $j$ draft tokens are accepted simultaneously [37], which is:
$AR_{i,j}=\prod_{k=0}^{j}p_{i,k}\approx\prod_{k=0}^{j}c_{i,k}$ (8)
Here, $p_{i,k}$ is the conditional probability of the $k^{th}$ draft token of the $i^{th}$ request is accepted given that all previous draft tokens have been accepted, and $c_{i,k}$ is its confidence score. Notably, we define $p_{i,0}=c_{i,0}=1$ to unify autoregressive decoding and speculative decoding.
The effectiveness of acceptance rate. Based on Eq. 8, the acceptance rate $AR$ exhibits two mathematical properties:
  1. 1.
    $AR\in[0,1]$, consistent with the nature of probability.
  2. 2.
    $AR_{i,j+1}\leq AR_{i,j}$, because $AR_{i,j+1}=AR_{i,j}*p_{i,j+1}$, and $p_{i,j+1}\leq 1$
At the same time, if the acceptance or rejection of draft tokens during the verification phase corresponds to an acceptance rate of 1 or 0, respectively, then the actual acceptance situation also conforms to the properties modeled by the acceptance rate, demonstrating the effectiveness of acceptance rate modeling in depicting actual acceptance situations.
$NAT=BS*\sum_{j=0}^{SL}\overline{AR_{j}}=BS*\overline{NAT}$ (9)
Further, as shown in Eq. 9, the mathematical properties of the average number of accepted tokens per request $\overline{NAT}$ can be derived as follows:
  1. 1.
    When $SL=0$, $\overline{NAT}=1$.
  2. 2.
    When $SL$ increases, $\Delta\overline{NAT}=\overline{AR}\in[0,1]$
  3. 3.
    When $SL$ increases, $\Delta\overline{NAT}$ is non-increasing.
Fig. 8 confirms the mathematical properties of the accepted token numbers and shows that the difference between the estimated number of accepted tokens and the actual number of accepted tokens is minimal, proving the effectiveness of the confidence-based method for estimating the number of accepted tokens.
4.3 The Relationship between Efficiency and Speculative Length
To achieve optimal efficiency through fine-grained control of speculative length, it is first necessary to understand the relationship between efficiency and speculative length, that is, how throughput varies with speculative length.
The relationship between execution time and speculative length. Given the average number of context tokens per request $\overline{N_{c}}$ and the batch size $BS$, we can derive $N_{dc}(i)$ and $N_{db}(i)$. Then, the speculative decoding time is given by:
$\begin{split}T_{sd}&=BS\big{(}\frac{SL^{2}(\alpha_{d}+\alpha_{v})}{2}\\ &\quad+SL(\alpha_{d}\overline{N_{c}}+\alpha_{v}\overline{N_{c}}-\frac{\alpha_{d}}{2}+\frac{\alpha_{v}}{2}+\gamma_{d}+\gamma_{v}+\frac{\delta_{d}}{BS})\\ &\quad+\alpha_{v}\overline{N_{c}}+\gamma_{v}+\frac{\delta_{v}}{BS}\big{)}\end{split}$ (10)
It is evident that the speculative decoding time is a quadratic function of the speculative length $SL$.
The relationship between throughput and speculative length. Throughput can be expressed as a function of speculative length, denoted as $f(s)=g(s)/h(s)$. Here, $g(s)$ represents the average number of tokens received per request ($\overline{NAT}$), and $h(s)$ represents the average speculative decoding time per request ($T_{sd}/{BS}$), where $s$ is the continuous speculative length. Herein, $g(s)$ and $h(s)$ possess the following properties:
  1. 1.
    $g(0)=1$, $0\leq g^{\prime}(s)\leq 1$, $g^{\prime\prime}(s)\leq 0$
  2. 2.
    $h(s)=as^{2}+bs+c>0$, $h^{\prime}(s)=2as+b>0$, $h^{\prime\prime}(s)=c>0$
The first derivative $f^{\prime}(s)$ and the second derivative $f^{\prime\prime}(s)$ of $f(s)$ can be represented as follows:
$f^{\prime}(s)=\frac{g^{\prime}(s)h(s)-g(s)h^{\prime}(s)}{\big{(}h(s)\big{)}^{2}}$ (11)
$\begin{split}f^{\prime\prime}(s)=&\frac{1}{\big{(}h(s)\big{)}^{4}}\Big{(}\big{(}g^{\prime\prime}(s)h(s)-g(s)h^{\prime\prime}(s)\big{)}\big{(}h(s)\big{)}^{2}\\ &-2h(s)h^{\prime}(s)\big{(}g^{\prime}(s)h(s)-g(s)h^{\prime}(s)\big{)}\Big{)}\end{split}$ (12)
It is evident that when $f^{\prime}(s)>0$, $f^{\prime\prime}(s)<0$. This characteristic indicates that as $s$ increases, $f^{\prime}(s)$ will eventually change within the negative range. Moreover, when $f^{\prime}(0)>0$, that is, when $g^{\prime}(0)c-b>0$, $f(s)$ initially increases and then decreases; otherwise, $f(s)$ monotonically decreases.
Furthermore, the variation of $f(s)$ is determined by:
$\begin{split}g^{\prime}(0)c-b=&2\alpha_{v}\overline{N_{c}}g^{\prime}(0)+2\gamma_{v}g^{\prime}(0)+\frac{2\delta_{v}}{BS}g^{\prime}(0)-2\alpha_{v}\overline{N_{c}}\\ &-\alpha_{v}-2\gamma_{v}-2\alpha_{d}\overline{N_{c}}+\alpha_{d}-2\gamma_{d}-\frac{2\delta_{d}}{BS}\end{split}$ (13)
, implying that throughput tends to first increase and then decrease over speculative length more readily under conditions including higher acceptance rates with greater $g^{\prime}(0)$, smaller batch sizes with lower $BS$, shorter contexts with fewer $\overline{N_{c}}$, larger target models with higher values of $(\alpha_{v},\gamma_{v},\delta_{v})$, and smaller draft models with lower values of $(\alpha_{d},\gamma_{d},\delta_{d})$.
In summary, we prove that there exists a unique optimal speculative length that can maximize the efficiency of speculative decoding. Therefore, we do not need to preset a range and iterate over all possible lengths to calculate the efficiency of speculative decoding; we only need to find this unique peak value.
5 Adaptive Speculative Decoding
Based on the relationship between efficiency and speculative length (§4.3), we propose adaptive speculative decoding. To implement this framework, we developed three key components. The first is the Adaptive Drafter (§5.1), which incorporates efficiency estimation into the drafting phase, achieving forward-looking step-level speculative length control. The second is the Confidence Prior Verifier (§5.2), which prioritizes the verification of tokens with high acceptance rates, achieving fine-grained request-level speculative length control. Lastly, the SLO-Aware Efficiency Estimator (§5.3) evaluates the efficiency of speculative decoding and achieves SLO awareness.
5.1 Adaptive Drafter
Algorithm 1 Adaptive Drafter
0:  Draft model $M_{d}$, Request set $reqs$, Confidence score of previous steps $prev\_conf$
0:  Draft token set $dtss$, Acceptance rate set $arss$
1:  $best\_thpt\leftarrow-1$, $dtss\leftarrow\{\}$, $arss\leftarrow\{\}$
2:  $prev\_confs\leftarrow\text{list}(prev\_conf)\times\text{lens}(reqs)$
3:  $arss.\text{append}(prev\_confs)$
4:  while True do
5:     $thpt\leftarrow\text{throughput\_estimation}(reqs,arss)$
6:     if $thpt>best\_thpt$ then
7:        $dts,confs\leftarrow M_{d}(reqs,dtss)$
8:        $dtss.\text{append}(dts)$
9:        $arss.\text{pop}()$
10:        if ${}arss==\{\}$ then
11:           $arss.\text{append}(confs)$
12:        else
13:           for $i\leftarrow 1$ to $\text{lens}(reqs)$ do
14:              $ars[i]\leftarrow arss[\text{lens}(arss)-1][i]\times confs[i]$
15:           end for
16:           $arss.\text{append}(ars)$
17:        end if
18:        $best\_thpt\leftarrow\text{throughput\_estimation}(reqs,arss)$
19:        $arss.\text{append}(prev\_confs)$
20:     else
21:        $arss.\text{pop}()$
22:        break
23:     end if
24:  end while
Based on the relationship between efficiency and speculative length, we have developed a module named adaptive drafter. This module is capable of adaptively adjusting the number of executions of the draft model during the drafting phase, in response to changes in the generated information. To achieve this goal, we designed a speculative controller that estimates its acceptance rate by analyzing the confidence score of draft tokens and decides whether to continue executing the draft model based on changes in the efficiency of speculative decoding. If it is found that continuing execution would reduce the efficiency, the controller will halt further execution of the draft model. However, the challenge with this approach is that the speculative controller needs to estimate the efficiency after execution based on the generated confidence scores, leading to the necessity of an additional execution of the draft model to ensure optimal efficiency is achieved.
To address this challenge, we adopted a "predict-execute-correct" strategy. As demonstrated in Alg. 1, we first use the average confidence scores from historical data to estimate the confidence scores that the next step of the draft model might produce, and based on this, predict the efficiency of speculative decoding. If the estimation indicates that continuing the execution of the draft model would improve the efficiency, the speculative controller executes the model, taking into consideration the newly generated draft tokens and their confidence scores, while also updating the collection of acceptance rates and the efficiency. If the estimation suggests that it is not appropriate to continue execution, the drafting phase is stopped. Through this approach, we are able to enable the draft model to rapidly generate draft tokens that are more likely to be accepted.
Algorithm 2 Confidence Prior Verifier
0:  Target model $M_{t}$, Request set $reqs$, Draft token set $dtss$, Acceptance rate set $arss$
0:  Output token set $otss$
1:  $best\_thpt\leftarrow\text{throughput\_estimation}(reqs,arss)$
2:  while True do
3:     $min\_ar,min\_ar\_token\leftarrow\text{find\_min\_ar}(dtss,arss)$
4:     $arss.\text{remove}(min\_ar)$
5:     $thpt\leftarrow\text{throughput\_estimation}(reqs,arss)$
6:     if $thpt>best\_thpt$ then
7:        $best\_thpt\leftarrow thpt$
8:        $dtss.\text{remove}(min\_ar\_token)$
9:     else
10:        break
11:     end if
12:  end while
13:  $otss\leftarrow M_{t}(reqs,dtss)$
5.2 Confidence Prior Verifier
We developed a module named the confidence prior verifier, which introduces an innovative request-level speculative length control strategy by decoupling the drafting and verification phases of speculative decoding. Currently, speculative length control is typically implemented on step-level [22], which limits the ability to adjust speculative lengths for different requests within the same batch individually. The confidence prior verifier achieves request-level speculative length control indirectly by prioritizing the verification of draft tokens anticipated to have a higher acceptance rate, thereby validating different numbers of draft tokens for different requests. Considering that the verification phase may occupy up to 90% of the total time for speculative decoding, this differentiated verification strategy among requests efficiently realizes fine-grained speculative length control.
As illustrated in Alg. 2, the confidence prior verifier employs a draft token elimination method to enhance the effectiveness of the draft token set entering the verification phase. This process starts with the collection of draft tokens and their estimated acceptance rates, evaluating whether a draft token should be eliminated in ascending order of acceptance rates. If the elimination operation can improve the efficiency of speculative decoding, the system executes the elimination and updates the collection of acceptance rates. This process is repeated until further elimination no longer enhances the efficiency. Ultimately, the Confidence Prior Verifier leverages the target model to validate the filtered draft tokens, achieving differentiated speculative decoding for different requests.
5.3 SLO-Aware Efficiency Estimator
Algorithm 3 SLO-Aware Efficiency Estimator
0:  Request set $reqs$, Acceptance rate set $arss$, SLO’s TPOT constraint $tpot$
0:  Estimated goodput $gdpt$
1:  $t\leftarrow\text{estimate\_time}(reqs)$
2:  if $t>tpot$ then
3:     $gdpt\leftarrow-1$
4:  else
5:     $nat\leftarrow\text{sum}(arss)$
6:     $gdpt\leftarrow\frac{nat}{t}$
7:  end if
To support the adaptive drafter and the confidence prior verifier, we design a SLO-aware efficiency estimator (Alg. 3). This estimator is capable of estimating the efficiency of speculative decoding based on the current workload and the speculative length.
Drawing from the model of accepted token numbers presented in §4.2, the SLO-aware efficiency estimator utilizes the acceptance rate predicted by confidence to determine the number of tokens to be accepted at the current step. Subsequently, in conjunction with the speculative decoding time model outlined in §4.1, it calculates the time required for speculative decoding based on the current workload. Through these two parameters, the estimator is able to estimate the speculative decoding throughput at the current step, thereby assisting SpecServe in achieving the optimal efficiency of speculative decoding.
Furthermore, we guide SpecServe in selecting the appropriate speculative length according to SLO requirements by setting constraints. Considering that in a batch of requests, some may only generate a single accepted token, making their TPOT equivalent to the time of speculative decoding. Therefore, to ensure that the time for each step of speculative decoding does not exceed the TPOT constraint specified by the SLO, strict control is necessary. If the time for speculative decoding exceeds the SLO limit, the throughput will be set to -1 to prevent the adaptive drafter from proceeding. In this manner, SpecServe can adjust the speculative length based on SLO constraints to better meet the SLO.
6 Implementation
SpecServe is implemented by extending the vLLM [17] with approximately 2000 lines of Python code. vLLM is a well-known production-level serving system that efficiently manages memory through PagedAttention and integrates online inference techniques such as continuous batching. At the speculative decoding mode, the lookahead scheduler within the vLLM divides the memory into two equal parts, allocating one for the KV Cache of the draft model and the other for the KV Cache of the target model. When a new batch of requests comes in, the adaptive drafter performs the draft model autoregressively through an "predict-execute-correct" process, maintaining a collection of acceptance rates until generating a set of draft tokens with the optimal efficiency of speculative decoding at the step level. Following this, the confidence prior verifier eliminates some draft tokens with low acceptance rates that hinder the efficiency of speculative decoding, and then inputs the set of draft tokens with the optimal efficiency at the request level into the target model for verification. After generating the results of the speculative decoding for this step, these results are integrated into the input of the speculative decoding for the next step.
Furthermore, to model the speculative decoding time, we introduce an offline analyzer. After each change of the draft and target model pairs and computing platform, the offline analyzer analyzes the forward pass time of the draft and target models under different batch sizes and context lengths, using a linear regression model as §4.1 to derive performance coefficients. During subsequent operations, these performance coefficients, combined with the batch size and context length processed in each step of speculative decoding, are used to estimate the time for each step of speculative decoding.
7 Evaluation
7.1 Experimental Setup
Table 1: Model pairs and hardware.
Target Model Draft Model Hardware
Vicuna-7B-v1.5 Vicuna-68M 1×L40
1×A800
Vicuna-33B-v1.3 Vicuna-160M 2×A800
4×A800
Llama-3.1-70B-Instruct Llama-3.2-1B-Instruct 8×A800
Model pairs and computing platforms. We conducted tests on three model pairs, which were completed across five different computational platforms. Specifically, we selected three target models: Vicuna-7B-v1.5 [43], Vicuna-33b-v1.3, and Llama-3.1-70B-Instruct [9], and matched them with Vicuna-68M [36], Vicuna-160M, and Llama-3.2-1B-Instruct as their respective draft models. Considering the memory resources, the model pair for the 7B target model operated on a single L40 or A800 GPU, the model pair for the 33B target model operates on either 2 $\times$ A800 GPUs or 4 $\times$ A800 GPUs, while the model pair for the 70B target model runs on 8 $\times$ A800 GPUs. In all these setups, the draft model operates with tensor parallelism set to 1, and only the target model is shared across multiple GPUs. We provide detailed specifications of the setup in Table 1.
Workloads. To evaluate the performance of SpecServe in real-world online inference systems, we selected three real-world production traces from the AzureLLMInferenceDataset2023 [29], specifically from the code and conversation categories, as well as mooncake [30], to simulate different request arrival patterns. To maintain system load within a normal range, we chose 10-minute segments with moderate request rates and adjusted the request rate of mooncake to one-third of its original rate. Consequently, these three traces respectively simulated scenarios of bursty requests (trace 1), high request rates (trace 2), and low request rates (trace 3).
To simulate the diversity of request types in real online inference systems, we created a mixed request set based on SpecBench [35], incorporating a variety of request types such as translation, summarization, question answering, mathematical reasoning, and retrieval-augmented generation. This request set was composed of 480 randomly selected instances from six widely used databases (MT-bench [43], WMT14 DE-EN, CNN/Daily Mail [27], Natural Questions [16], GSM8K [7], and DPR [14]).
Figure 9: Speedup in end-to-end latency across different traces, model pairs and computing platforms.
Figure 10: SLO attainment and speedup in end-to-end latency across different traces, model pairs and computing platforms.
Baselines. We compared SpecServe with four baseline methods: (1) autoregressive decoding; (2) vanilla speculative decoding with preset speculative lengths of 1, 3, and 5; (3) threshold-based dynamic speculative decoding, a dynamic speculative decoding method that has been integrated into the transformer library [34], representing the state-of-the-practice method; (4) SpecInfer [26], a state-of-the-art speculative decoding method integrated into inference systems. Due to the inefficiency of tree-based inference strategies (with a tree width of 3, the acceptance rate of draft tokens does not exceed 18% [26]), we set the tree width to 1 and name it SpecInfer*.
Metrics. To evaluate the reduction in latency, we measure the end-to-end latency speedup relative to autoregressive decoding. Drawing upon prior research [44], we established the TTFT at 200ms and the TPOT at 30ms as our SLO metrics. By adjusting the scale factor within the range of 0.8 to 1.4, we simulated varying degrees of SLO stringency.
7.2 Overall Performance of SpecServe
As illustrated in Fig. 9-11, we first evaluated the end-to-end latency speedup and SLO attainment of all methods under real-world production traces, and assessed the end-to-end latency speedup of all methods across various batch sizes.
Consistent low inference latency. As shown in Fig. 9, we observe that SpecServe’s speedup ranged from 1.1× to 1.8× relative to autoregressive decoding across 15 experimental setups, demonstrating its stability in reducing latency. In contrast, vanilla speculative decoding methods (with speculative lengths of 1, 3, 5) have a speedup ratio of less than 1 in 7%, 13%, and 40% of the experimental setups, respectively, indicating that maintaining low latency becomes increasingly challenging with longer preset lengths. Threshold-based dynamic method does not achieve the expected outcome in 40% of the setups, suggesting that a single confidence score is insufficient for accurately determining the optimal speculative length. SpecInfer* is the least stable, primarily due to its inefficient speculative strategy. By adapting to different request patterns and system configurations, SpecServe consistently optimized end-to-end latency, making it a reliable inference engine suitable for real-world production environments.
Compared with the optimal methods in various scenarios, SpecServe outperforms the baseline method in terms of latency in 10 out of 15 experimental setups, and matches the latency performance of the baseline method in the remaining 5 setups. From a system configuration perspective, SpecServe demonstrates a more significant latency reduction in resource-constrained scenarios, such as the deployment of a 33B+160M model pair on a 2×A800 computing platform compared to the same model pair on a 4×A800 computing platform. From the perspective of workload, when facing scenarios with bursty requests (trace 1), SpecServe is capable of achieving an additional speedup of 5% to 10% compared to other baseline methods. This indicates that SpecServe’s fine-grained speculative length control method can adaptively determine the optimal speculative length, especially suitable for resource-limited and complex request pattern production environments.
Figure 11: Speedup in end-to-end latency across various batch sizes.
High SLO attainment. Fig. 10 demonstrates the SLO attainment and speedup of different methods under request burst scenarios (trace 1) with varying SLO constraints. As depicted, SpecServe’s SLO attainment rate exceeds that of other methods by 5% to 38% across all SLO constraints, and its speedup relative to autoregressive decoding surpasses other methods by 10% to 61%. Furthermore, SpecServe can adjust its speculative length according to different SLO constraints, thereby altering its speedup. For instance, under loose SLO constraints, SpecServe can achieve an additional 2% speedup. This indicates that SpecServe not only attains a higher speedup while maintaining or exceeding the SLO attainment of autoregressive decoding but also flexibly adjusts according to varying SLO constraints to optimize performance. This characteristic effectively mitigates the issue of speculative decoding, where pursuing a higher speedup often results in a lower SLO attainment, making SpecServe a fast and SLO-aware solution.
Batch adaptivity. As shown in Fig. 11, under arbitrary batch sizes composed of different requests, SpecServe achieves the highest speedup among all compared baseline methods. Its superior batch adaptability accounts for the demonstrated stable latency reduction and low-latency characteristics. In contrast, vanilla speculative decoding with a preset length cannot match the batch adaptability of dynamic methods. These static method either employ conservative speculative lengths to maintain a relatively stable reduction in latency or use aggressive speculative lengths to gamble for a higher speedup, unable to achieve both simultaneously. Obviously, dynamic methods also require a reasonable speculative length control scheme. Threshold-based approaches can only offer random improvements in speedup and are incapable of maintaining stable latency reduction. In summary, SpecServe’s superior batch adaptability is particularly crucial for production environments, as it ensures that SpecServe can maintain efficient batch processing capabilities most of the time, even in the face of significant fluctuations in batch size.
7.3 Ablation Study
Table 2: The variations in average speculative length, average number of accepted draft tokens, average draft token acceptance rate, and speedup across different traces for various dynamic methods.
Trace Method # Draft Tokens (Avg.) # Accepted Draft Tokens (Avg.) Acceptance Rate SpeedUp
Threshold-based 6.54 1.49 0.20 0.54×
Trace 1 SpecServe- 1.64 0.63 0.40 1.09×
SpecServe 1.59 0.63 0.42 1.13×
Threshold-based 5.67 1.44 0.20 0.85×
Trace 2 SpecServe- 3.26 1.18 0.39 1.34×
SpecServe 3.17 1.16 0.40 1.38×
Threshold-based 3.84 1.34 0.31 1.51×
Trace 3 SpecServe- 4.20 1.34 0.33 1.57×
SpecServe 4.08 1.33 0.34 1.58×
As shown in Table 2, we conducted an ablation study of the performance of different dynamic methods on real-world production traces, exploring the sources of speedup for each method across three dimensions: average speculative length, average accepted draft tokens, and average draft token acceptance rate. Specifically, SpecServe- solely utilizes an adaptive drafter, without employing the confidence prior verifier.
Benefits of adaptive drafter. Effective control of speculative length has achieved an additional acceleration ratio of 6% to 55%. We observe that SpecServe- can achieve an average draft token acceptance rate that is twice as high as that based on the threshold-based method, and the higher the acceptance rate, the greater the additional speedup achieved by SpecServe-. This indicates that the speedup of SpecServe- primarily benefits from its ability to efficiently convert draft tokens into accepted tokens. Crucially, the adaptive drafter enhances the average draft token acceptance rate by adopting longer speculative lengths for batches that are likely to accept more tokens. Unlike the threshold-based method that only control speculative length by comparing confidence scores, the adaptive drafter employs a model based on confidence scores to estimate the expected number of accepted tokens per batch, achieving more precise control of speculative length.
Benefits of confidence prior verifier. Request-level control of speculative length further enhances the acceleration ratio by an additional 2% to 5%. SpecServe possesses the highest average draft token acceptance rate among all dynamic methods, which is precisely why it achieves the highest speedup. Compared to SpecServe-, SpecServe slightly affects the average number of accepted draft tokens (a reduction of only 10% of the average speculative length) while reducing the average speculative length, resulting in an increase in the average draft token acceptance rate by 2% to 4%. This indicates that for every ten draft tokens eliminated by the confidence prior verifier, only one could have been accepted, demonstrating the effectiveness of the draft token elimination strategy.
7.4 Behavior Analysis
Figure 12: Fluctuations in batch sizes and speculative lengths under different methods and traces.
As illustrated in Fig. 12, we analyzed the behavior of optimal vanilla speculative decoding (with a speculative length of 3), threshold-based dynamic speculative decoding, and SpecServe across traces from three real production environments. Within a 10-minute observation period, we monitored the current batch size and speculative length every 5 seconds.
Effective speculative length control. Across different production traces, SpecServe’s speculative length generally exhibits an inverse relationship with batch size changes, i.e., as batch size increases, speculative length decreases. In contrast, the speculative length of threshold-based dynamic speculative decoding is unrelated to batch size, often utilizing the maximum speculative length of 8, resulting in a higher average speculative length, as observed in §7.3. On the other hand, the preset length of vanilla speculative decoding cannot adjust speculative length, and its experimental outcomes rely entirely on empirical data. Therefore, through effective and fine-grained speculative length control, SpecServe can adapt to varying request patterns and system configurations, making it highly suitable for integration into actual production environment service systems.
System behaviors. SpecServe generally exhibits lower batch size peaks in behavior analysis, indicating superior performance in handling system request backlogs. Batch size essentially reflects the system’s request backlog volume. Ideally, the system aims to maintain a constant request backlog volume, meaning the number of requests processed by the system within a specific timeframe equals the number of requests received, without resource wastage or shortage. However, in real-world scenarios, especially during sudden high request rates (e.g., the high request rate burst starting at the 410-th second in trace 1), the request backlog volume may sharply increase, placing pressure on system operation. In such instances, SpecServe’s batch size peak is only 43% of that achieved by the threshold-based method and 88% of the optimal vanilla speculative decoding. This effectiveness is primarily attributed to SpecServe’s effective speculative length control mechanism. Faced with a large volume of requests, SpecServe can automatically switch to autoregressive decoding mode, maintaining lower end-to-end latency and effectively avoiding request backlogs. Overall, SpecServe’s stable latency reduction feature achieves smoother system operation, effectively addressing sudden request spikes in real scenarios.
8 Related Work
Speculative decoding solutions. The efficiency of speculative decoding is influenced not only by the speculative length but also depends on the execution efficiency of the draft model and the probability of the draft tokens being accepted by the target model. Hence, developing draft models that are both efficient and similar to the target model has become a focus of speculative decoding research. Previous studies have mainly adopted two approaches to designing draft models: (1) Independent draft models, which involve introducing a high-efficiency independent draft model to generate draft tokens. This model could be a variant from the same series as the target model [18, 33], a fine-tuned model [26, 15, 23, 45], or based on databases from other textual resources [13]; (2) Self-draft models, which adjust the structure of the target model to enable it to rapidly generate multiple draft tokens [4, 19, 11, 42, 20]. Moreover, considerable work has applied speculative decoding to specific scenarios, such as tasks involving small batches of ultra-long sequences [32] and large batches of medium-length sequences [6], designing specific efficient draft models. The diversity of draft models has led to the flourishing development of the speculative decoding, and SpecServe can adaptively control the speculative length for any type of draft model with fine granularity.
AdaServe [21] and our work are contemporaneous, with the following differences: (1)AdaServe focuses on providing SLO-customized services for requests, while SpecServe is committed to achieving stable acceleration guarantees for LLM inference under dynamic scenarios and strict SLO constraints. (2)AdaServe only requires the expected value of TPOT to meet the SLO, whereas SpecServe proposes that the worst-case scenario of TPOT (i.e., the upper limit of the speculative decoding step time) must always strictly comply with the SLO constraints. (3)AdaServe primarily targets tree-based speculative decoding, while SpecServe specifically optimizes for sequential speculative decoding.
Resource-efficient LLM serving. In addition to speculative decoding, existing LLM serving systems propose other solutions to improve resource efficiency. For instance, Orca [38] leverages continuous batching to reduce request queuing delays, thereby achieving efficient GPU resource utilization; vLLM [17] constructs a memory-efficient service system based on PagedAttention; while Llumnix [31] focuses on multi-tenant scenarios, enabling efficient cross-node resource utilization through runtime rescheduling across multiple model instances. Furthermore, the disaggregated architecture [29, 30, 44] has also become a focal point, where the disaggregation of prefill and decoding The separation of prefill and decoding can optimize resource utilization and achieve a balance between TTFT and TPOT. SpecServe enables speculative decoding to serve as an efficient and stable decoding solution, forming an orthogonal synergy with the resource-efficient techniques of the aforementioned serving systems.
9 Conclusion
Speculative decoding aims to reduce inference latency by increasing computational overhead, yet this trade-off depends on the variations in request patterns, system configurations, and SLOs. We explore how these variables influence speculative decoding and have developed SpecServe, an adaptive speculative decoding framework that addresses these variations through request-level speculative length control. Experimental results demonstrate that SpecServe can effectively and consistently reduce inference latency while maintaining lower computational costs.